OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 − 6x + 2 in vertex form.

OpenStudy (anonymous):

f(x) = x² - 6x + 2 f(x) - 2 = x² - 6x f(x) - 2 + 9 = x² - 6x + 9 f(x) + 7 = (x - 3)² f(x) = (x - 3)² - 7 Could you please fan me?

OpenStudy (danjs):

hah

OpenStudy (solomonzelman):

this is against the rules, and should be agains your conscience as well (if you got one)

OpenStudy (danjs):

don't just give solutions, help the person along prolly

OpenStudy (anonymous):

thank you anyway tho aha

OpenStudy (anonymous):

i didn't !

OpenStudy (anonymous):

f(x) = x² - 6x + 2 f(x) - 2 = x² - 6x f(x) - 2 + 9 = x² - 6x + 9 f(x) + 7 = (x - 3)² f(x) = (x - 3)² - 7 Those are the steps!